40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Apr 13 03:15:12 2012, koji,frank,tara, DailyProgress, optic, vibration damping on optics IMG_0731.jpgdamped.png
    Reply  Sat Apr 14 06:03:23 2012, frank,tara, DailyProgress, optic, vibration damping on optics IMG_0744.jpgIMG_0742.jpg
       Reply  Mon Apr 16 02:20:58 2012, tara, DailyProgress, optic, vibration damping on optics beat_2012_04_15.pngbeat_2012_04_15.png
          Reply  Mon Apr 16 22:02:08 2012, tara, DailyProgress, optic, vibration damping on optics beat_2012_04_15_fixed.png
             Reply  Wed Apr 18 00:49:07 2012, tara, DailyProgress, optic, vibration damping on optics 
Message ID: 923     Entry time: Wed Apr 18 00:49:07 2012     In reply to: 921
Author: tara 
Type: DailyProgress 
Category: optic 
Subject: vibration damping on optics 

I planned to measure the beat at night with the air springs activated, but the power went out around 11:45 pm. I think the temperature servo got a kick and it is drifting very fast. So I cannot keep the cavities locked long enough for the low frequency measurement. I'm just turning the systems back on for now.

The laser, 3 Marconis for 14.75MHz EOM, for ACAV AOM, for beat are set back to the original setup, PMC medm screen are back on, the air springs are up and working.

The linux machine is on but I forgot the password, will ask Frank tomorrow.

ELOG V3.1.3-